OpenStudy (anonymous):

solve and check. 26x^2-9x=0

geerky42 (geerky42):

Factor x out: x(26x - 9) = 0 x = 0 26x - 9 = 0 26x = 9 x = 9/26 So x = 0 or 9/26 Now plug in 0 or 9/26 and check.

OpenStudy (anonymous):

thanks alot :) you made it easy to follow and understand, my answer was completely wrong lol. thanks for helping me

geerky42 (geerky42):

You're welcome.